# Scalable Multi-Robot Motion Planning Using Guidance-Informed Hypergraphs

### v3

@@ IV Workspace Guided-DaSH
In this section, we explain how we apply the Guided-DaSH framework for mobile robot MRMP in congested environments with narrow passages to propose the Workspace Guided-DaSH method and describe how we leverage the workspace skeleton to guide both the construction of $\mathcal{H}_{ts}$ and its translation into $\mathcal{H}_{m}$.
IV-A Task Space Hypergraph Construction
@@
Considering mobile robots, coordination between the motions of robots is required when they are physically in proximity to each other. Our insight is that we can leverage the skeleton to approximate when coordination is necessary by modeling the movement of the robots through the workspace as movement along the skeleton. Specifically, coordination is necessary when robots are traversing the same skeleton edge (either in the same or opposite direction) or traveling through the same skeleton vertex. This formulation imposes a structure on the search space that we capture through our heuristic search.
@@
Workspace Guided-DaSH follows the structure shown in Alg. 1. Given a skeleton of the workspace, we first conduct our heuristic search to find a path over the skeleton for each robot using a MAPF algorithm. We set the capacity of each skeleton vertex and edge as given by the minimum clearance to an obstacle. We then combine these paths to form the minimal necessary portion of the task space hypergraph. Fig. 2 shows an example of the conversion of the MAPF solution into $\mathcal{H}_{ts}$. Within task space elements, constraints represent the skeleton segments that each robot should traverse at that stage. Robots traversing the same skeleton edge are grouped into the same task space element using composition hyperarcs. As robots move away from each other onto different edges, they are decoupled into separate task space elements, boosting planning efficiency by considering smaller $\mathcal{C}_{spaces}$. Transition hyperarcs capture robots moving through the same skeleton vertex as well as robots moving between segments of the skeleton (i.e. moving between task space elements with the same robots but different constraints).
@@
Figure 2: A segment of the robot paths over the skeleton (gray) and the corresponding portion of the task space hypergraph. Each task space element shows the corresponding set of robots and the relevant portion of the topological skeleton. In (b), arrows indicate robots moving along a skeleton edge and dots indicate a decoupled state. Solid lines indicate transition hyperarcs and dashed lines represent composition hyperarcs. The green and red robots are initially moving in opposite directions along the same skeleton edge. Then, red and blue reach the skeleton vertex together and go on to different edges. The yellow robot begins moving in the opposite direction of the red robot when the red robot reaches the next skeleton edge.
IV-B Translation to Low-level Motion Hypergraph
@@
Once $\mathcal{H}_{ts}$ is constructed, we compute a set of paths to form $\mathcal{H}_{m}$. To ensure that the robots roughly follow their paths over the skeleton and that narrow passages in the environment are efficiently traversed, we leverage the skeleton to guide the construction of local paths along skeleton edges. To compute paths through task space elements, we use the dynamic sampling region advancement procedure from CDR-RRT [6].
@@
Considering transition hyperarcs that represent multiple robots moving through a skeleton vertex, we stop the movement of the robots along the incoming skeleton edges a small distance $\delta$ from the end of each edge. We’ve found that setting $\delta$ to the diameter of the region used for the dynamic sampling region advancement works well. We then grow an RRT [14] in the composite space of these robots from the last configuration on the robots’ incoming skeleton edges to a sampled configuration at least $\delta$ distance away from the skeleton vertex of the robots’ next skeleton edges. Here, we simply grow an RRT rather than using CDR-RRT to avoid forcing the robots to reach the skeleton vertex at the same time, creating congestion. Composition hyperarcs and transition hyperarcs not associated with movement in the task space hypergraph correspond directly to vertices and hyperarcs in the motion hypergraph.
@@
We consider construction of a local path to have failed if the RRT fails to be extended a given number of times during planning. In this case, we impose a constraint that the MAPF solution from which we build $\mathcal{H}_{ts}$ cannot contain the same group of robots traversing along either a skeleton edge that failed or the same groups of robots traversing along the incoming and outgoing skeleton edges to a skeleton vertex that failed. We then recompute $\mathcal{H}_{ts}$ and try again to construct $\mathcal{H}_{m}$. Similarly, if conflicts are found between robots traversing different portions of the skeleton simultaneously, constraints are imposed to prevent the relevant robot groups from traversing these segments of the skeleton together.
IV-C Kinodynamic Planning
@@
Incorporating kinodynamic constraints into Workspace Guided-DaSH is straightforward due to the modularity of the framework. As kinodynamic constraints influence the motion of the robots, it is imperative that the low-level motion hypergraph, $\mathcal{H}_{m}$, incorporates the kinodynamic constraints during construction. To this end, we adapt the use of CDR-RRT, employed during the construction of $\mathcal{H}_{m}$ to incorporate kinodynamic constraints. This is achieved by substituting their underlying RRT with Kinodynamic RRT [15].
@@
A control input is generated by sampling the control space uniformly at random which is then integrated forward during the RRT extension phase for a randomly sampled duration. In order to utilize the guidance provided by the CDR-RRT, we generate and integrate forward many random control inputs and select the input which moves the system closest to the dynamic sampling region. The kinodynamic RRT continues to expand until either a maximum number of states have been added or the goal region has been reached. We refer the readers to [15] for more details on kinodynamic RRT.
@@ V Validation
V-A Experimental Setup
We measure the performance of Workspace Guided-DaSH against CDR-RRT [6] as well as a set of other approaches that are relevant for comparison. Composite RRT [14] operates in the full composite space, providing a high level of coordination. MRdRRT [16] was originally intended for manipulators, but its implicit search of the composite space makes it pertinent for comparison. sRRT [19] uses subdimensional expansion to limit the size of the search space which is similar in concept to our hypergraph representation. ARC [20] composes robots into higher dimensional planning spaces based on encountered conflicts, whereas our method uses workspace guidance to preemptively avoid likely conflicts. Thus, comparing our method against ARC effectively measures the impact of this guidance. We use Conflict-based Search [25] over the workspace skeleton to perform the heuristic search for Workspace Guided-DaSH. Each method was implemented in the open-source Parasol Planning Library. Holonomic experiments were run on a 4-core CPU (4GHz, 32GB RAM), while kinodynamic experiments used a 24-core CPU (6GHz, 64GB RAM). We run 15 seeds for each scenario. Each method is given 600 seconds to solve a scenario or is considered a failure. We report planning time as well as path cost given by the makespan.
@@ V-A Experimental Setup
V-B Environments
Here, we describe our experimental environments and explain which aspects of our approach these scenarios are designed to highlight.
V-B1 Warehouse (Fig. 3(a))
The Warehouse environment features several narrow aisles through which robots must swap places with the robot opposite them. In the holonomic scenario, the width of the narrow passages is approximately 2.5 times the diameter of the disk robots we consider. We additionally run a kinodynamic variant with doubled aisle widths. This scenario measures each approach’s ability to efficiently handle narrow passages within the environment. We highlight Workspace Guided-DaSH’s ability to decompose the problem using topological guidance.
V-B2 Tunnels (Fig. 3(b))
The Tunnels environment again features several narrow passages. Here, because of the 3D workspace, the planning space becomes very large. This scenario is intended to measure each algorithm’s performance with respect to the geometric complexity, rather than the congestion, of the environment. We consider L-shaped robots with the width of their longest dimension slightly less than half the width of the narrow passages. Each robot must find and traverse multiple narrow passages to reach its goal, but its shortest path does not conflict with any other robot’s.
@@ V-B Environments
V-B3 Mining (Fig. 3(c))
The Mining environment is a 3D environment meant to mimic a set of mine shafts. Again, it features several narrow passages; however, here robots must move in close proximity to each other. We use the same robots as the Tunnels environment. We show that Workspace Guided-DaSH is capable of efficiently finding paths for large robot teams in problems representative of real-world scenarios.
V-B4 GridMaze (Fig. 3(d))
@@
V-C Experimental Results
In the Warehouse scenario (results in Tab. I), all methods except MRdRRT achieve a 100% success rate in the 2-robot scenario. However, the methods that do not leverage topological guidance begin to struggle with the 4-robot scenario and multiple cannot solve the 8-robot scenario. Of those, only ARC is able to solve up to the 32-robot scenario; however, with a success rate of only 26.6%. CDR-RRT and sRRT are able to solve up to the 8-robot scenario where the exponential size of the composite space leads to a significant decrease in performance. Workspace Guided-DaSH, like CDR-RRT, sees the benefit of using the skeleton representation to guide planning through narrow passages and the coordination provided by using the composite space of multiple robots. Additionally, by using the workspace to inform the planner when to move between different levels of composition, Workspace Guided-DaSH is able to focus planning within smaller spaces than CDR-RRT, making it able to plan for significantly larger robot teams, up to 128 robots, within the 600-second time limit.
In the Tunnels environment (results in Tab. II), the size of the search space is increased significantly relative to the Warehouse environment since the robot now has six degrees of freedom. Composite RRT, MRdRRT, and ARC are unable to find a solution for the 2-robot scenario due both to the large size of the search space and the complexity of the environment. sRRT performs better because of its ability to decompose the search space, but struggles with the narrow passages and only one seed was able to solve the 4-robot scenario. Here, the narrow passage problem is exacerbated since the query requires the planner to find and traverse multiple narrow passages. By leveraging skeleton guidance, CDR-RRT and Workspace Guided-DaSH are able to find solutions for larger groups of robots. Searching the composite space hinders CDR-RRT’s performance, leaving it unable to solve the 64-robot scenario within the time limit. Workspace Guided-DaSH is able to efficiently find solutions for up to 64 robots within the time limit due to its ability to decompose the planning space.
@@ V-C Experimental Results
In the Mining environment (results in Tab. III), Composite RRT and MRdRRT are unable to find a solution for the 2-robot scenario due to the large size of the search space and inter-robot conflicts. Similar to the Tunnels scenario, ARC’s performance is impacted by the environment complexity since it does not use topological guidance and it fails to find solutions for the 2-robot scenario within the time limit. sRRT struggles to find solutions for the 2-robot scenario and cannot solve the 4-robot scenario. By leveraging topological guidance to limit exploration of the composite space, CDR-RRT is able to solve up to the 16-robot scenario. However, it is unable to solve the 32-robot scenario. Workspace Guided-DaSH, by considering the composite space only when necessary, is able to efficiently solve up to the 64-robot scenario.
@@
In the 3D GridMaze scenario (results in Tab. IV), as a result of the large size of the planning space as well as the difficulty posed by the long curved narrow passages, Composite RRT struggles with the 2-robot scenario and fails to produce a solution for the 4-robot scenario. MRdRRT fails to solve the 2-robot scenario. While CDR-RRT is able to solve the 4-robot scenario, the large size of the composite space prevents it from finding a solution for the 8-robot scenario within the time limit. sRRT and ARC benefit from searching a lower-dimensional space when robots are not in conflict, but fail to solve the 32-robot scenario where conflicts become prevalent due to the high level of congestion of the environment. Similar to sRRT, Workspace Guided-DaSH searches lower-dimensional spaces, however, our method also uses the workspace skeleton for conflict resolution. When a conflict is too difficult to be resolved quickly via random sampling, we impose constraints on the robots’ abstract paths through the workspace which allows us to find conflict-free paths more efficiently. Thus, Workspace Guided-DaSH is able to find solutions up to 32 robots. Although this replanning mechanism sometimes results in slightly higher solution costs relative to other methods, a post-processing trajectory optimization step is generally used in practice to smooth paths.
V-D Kinodynamic Experiments

### v4

@@ III Workspace Guided-DaSH Method
In this section, we propose the Workspace Guided-DaSH method, which leverages the workspace skeleton to guide both the construction of $\mathcal{H}_{ts}$ and its translation into $\mathcal{H}_{m}$. Fig. 1 compares the DaSH framework with our method, which we elaborate on below.
III-A Problem Formulation
We consider the problem of multiple mobile robot motion planning in congested environments with narrow passages. Given a set of $n$ robots, a set of start and goal configurations, $q_{s}^{i}~\forall i\in[1,n]$ and $q_{g}^{i}~\forall i\in[1,n]$ respectively, we aim to find a path for each robot $\pi^{i}~\forall i\in[1,n]$ that satisfies the criteria that $\pi^{i}_{0}=q_{s}^{i}~\forall i\in[1,n]$ and $\pi^{i}_{T}=q_{g}^{i}~\forall i\in[1,n]$ where $\pi^{i}_{t}$ represents the path configuration at time $t$ and $T$ is the final timestep. Additionally, these paths must be inter-robot collision free. We consider two path configurations to be in collision if the geometry of the robots at these configurations intersect. Our planner is generic and we do not make assumptions about robot geometry or kinematics. We assume a known environment and precompute a workspace skeleton representation.
III-B High Level Task Space Hypergraph Construction
@@
Constructing $\mathcal{H}_{ts}$ requires knowledge about the problem space to make composition decisions about when coordination between robots is required during planning and the relevant path constraints. Considering mobile robots, coordination between robots’ motions is required when they are physically near each other. Our insight is that we can leverage the skeleton to approximate when coordination is needed by modeling the movement of the robots through the workspace as movement along the skeleton. Specifically, coordination is needed when robots are traversing the same skeleton edge (either in the same or opposite direction) or traveling through the same skeleton vertex. This formulation imposes a structure on the search space.
@@
Workspace Guided-DaSH follows the structure shown in Alg. 1. Given a workspace skeleton, we first conduct a heuristic search to find a path over the skeleton for each robot using a MAPF algorithm [ssfs-cbsfomap-15] (line 5). We set the capacity of each skeleton vertex and edge as given by the minimum clearance to an obstacle. We then combine these paths to form the minimal necessary portion of $\mathcal{H}_{ts}$ (line 6). Fig. 2 shows an example of the conversion of the MAPF solution into $\mathcal{H}_{ts}$. Within task space elements, constraints represent the skeleton segments that each robot must traverse at that stage. Robots traversing the same skeleton edge are grouped into the same task space element. As robots move away from each other to different edges, they are decoupled into separate task space elements, boosting planning efficiency by considering smaller $\mathcal{C}_{spaces}$.
@@
Figure 2: (a) A segment of the robot paths over the skeleton (gray) and (b) the corresponding portion of the task space hypergraph. Each task space element (circle) shows the corresponding set of robots and the relevant portion of the workspace skeleton. In (b), arrows indicate robots moving along a skeleton edge and dots indicate a decoupled intermediate state. The green and red robots are initially moving in opposite directions along the same skeleton edge. Then, red and blue reach the skeleton vertex together and go on to different edges. The yellow robot begins moving in the opposite direction of the red robot when the red robot reaches the next skeleton edge.
III-C Translation to Low-level Motion Hypergraph
@@
The motion hypergraph $\mathcal{H}_{m}$ stores the configurations and motions between them that make up the robots’ paths through the task space elements and hyperarcs from $\mathcal{H}_{ts}$. For each task space element and hyperarc $x_{ts}\in\mathcal{V}_{ts}\cup\mathcal{E}_{ts}$, ordered from start to goal (line 7), a path is computed in the composite space of the robots contained therein (line 10). These paths become vertices and hyperarcs in $\mathcal{H}_{m}$. To ensure that the robots roughly follow their paths over the skeleton and that narrow passages in the environment are efficiently traversed, we leverage skeleton guidance during path construction. To compute paths through task space elements, we use the dynamic sampling region advancement procedure from CDR-RRT [mmuma-smrmpcetg-23].
@@
Considering hyperarcs that represent multiple robots moving through a skeleton vertex, we stop the movement of the robots along the incoming skeleton edges a small distance $\delta$ from the end of each edge to avoid conflicts while moving into the vertex. We’ve found that setting $\delta$ to the diameter of the region used for the dynamic sampling region advancement works well. We then grow an RRT [l-rrtntpp-1998] in the composite space of these robots from the last configuration on the robots’ incoming skeleton edges to a sampled configuration at least $\delta$ distance away from the skeleton vertex of the robots’ next skeleton edges. Here, we simply grow an RRT rather than using CDR-RRT to avoid forcing the robots to reach the skeleton vertex at the same time, creating congestion.
@@
We consider construction of a path to have failed if the RRT fails to be extended a given number of times during planning (line 12). In this case, we impose a constraint that the MAPF solution from which we build $\mathcal{H}_{ts}$ cannot contain the same group of robots traversing along either a skeleton edge that failed or the same groups of robots traversing along the incoming or outgoing skeleton edges to a skeleton vertex that failed. We then recompute $\mathcal{H}_{ts}$ (line 6) and try again to construct $\mathcal{H}_{m}$.
Since we know from the heuristic search over the skeleton which task space elements’ motions occur concurrently, we can interleave motion hypergraph construction and conflict detection. By finding conflicts between different task space elements as they occur, we minimize wasted computation and can directly extract the motion solution from the motion hypergraph, bypassing DaSH’s query stage. As we construct $\mathcal{H}_{m}$, after each path is computed, we check for collisions with existing paths occurring simultaneously by considering intermediate configurations along the paths (line 14). If conflicts are found between robots traversing different portions of the skeleton simultaneously, similarly to the constraints added when path construction fails, MAPF constraints are imposed to prevent the relevant robot groups from traversing these segments of the skeleton together (line 16).
Our modification to the conflict resolution stage also allows us to support kinodynamic planning. The scheduled query proposed in the original DaSH framework resolves conflicts by inducing waiting to prevent robots from colliding with each other. When planning under kinodynamic constraints, we cannot assume the ability of robots to instantaneously change velocity. Additionally, to avoid solving computationally difficult optimal control point-to-point problems, we cannot reuse paths for task space elements occurring after conflicts in the robots’ paths. Thus, we must replan affected paths to avoid conflicts.
After new MAPF constraints are added, the heuristic search recomputes $\mathcal{H}_{ts}$ (line 6), and construction of $\mathcal{H}_{m}$ restarts. This process continues until a conflict-free motion solution is found (lines 1819).
III-D Kinodynamic Planning
@@
As kinodynamic constraints influence the motion of the robots, it is imperative that the low-level motion hypergraph, $\mathcal{H}_{m}$, incorporates the kinodynamic constraints during construction. This is achieved by substituting the underlying RRT with Kinodynamic RRT [hklr-rkmpwmo-02].
@@
A control input is generated by sampling the control space uniformly at random, which is then integrated forward during the RRT extension phase for a randomly sampled duration. In order to utilize the guidance provided by the CDR-RRT, we generate and integrate forward many random control inputs and select the input that moves the system closest to the dynamic sampling region. The kinodynamic RRT continues to expand until either a maximum number of states have been added, a failure, or the goal region has been reached.
@@ IV Validation
IV-A Experimental Setup
We measure the performance of Workspace Guided-DaSH against CDR-RRT [mmuma-smrmpcetg-23] as well as a set of other relevant approaches. Composite RRT [l-rrtntpp-1998] operates in the full composite space, providing high coordination. MRdRRT [ssh-faniaehdrfeoirimm-16] was originally intended for manipulators, but its implicit search of the composite space makes it pertinent for comparison. sRRT [wkc-pppfmrwse-12] uses subdimensional expansion to limit the size of the search space which is similar in concept to our hypergraph representation. ARC [smqma-arc-24] composes robots into higher-dimensional planning spaces based on encountered conflicts, whereas our method uses workspace guidance to preemptively avoid likely conflicts. Thus, comparing our method against ARC effectively measures the impact of this guidance. Each method was implemented in an open-source library and code will be provided in the final manuscript. In all environments, we find plans for holonomic robots considering only geometric feasibility, deferring kinodynamic feasibility to a post-processing trajectory optimization step. In the Warehouse environment, we additionally incorporate kinodynamic constraints into planning. We run 15 seeds for each scenario. Each method is given 600 seconds to solve a scenario or is considered a failure. We report planning time as well as path cost given by the makespan.
@@ IV-A Experimental Setup
IV-B Environments
Here, we describe our experimental environments and explain which aspects of our approach they highlight.
IV-B1 Warehouse (Fig. 3(a))
The Warehouse environment features narrow aisles through which the robots must swap places. In the holonomic scenario, the width of the narrow passages is 2.5 times the diameter of the disk robots we consider. We additionally run a kinodynamic variant with doubled aisle widths. This scenario measures each approach’s ability to efficiently handle narrow passages within the environment. We highlight Workspace Guided-DaSH’s ability to decompose the problem using topological guidance.
IV-B2 Tunnels (Fig. 3(b))
The Tunnels environment again features narrow passages. Because of the 3D workspace, the planning space becomes very large. This scenario is intended to measure each algorithm’s performance with respect to the geometric complexity, rather than the congestion, of the environment. We consider L-shaped robots with the width of their longest dimension slightly less than half the width of the narrow passages. Each robot must find and traverse multiple narrow passages to reach its goal, but its shortest path does not conflict with any other robot’s.
@@ IV-B Environments
IV-B3 Mining (Fig. 3(c))
The Mining environment is a 3D environment meant to mimic a set of mine shafts. Again, it features narrow passages; however, here robots must move in close proximity to each other. We use the same robots as the Tunnels environment. We show that Workspace Guided-DaSH is capable of efficiently finding paths for large robot teams in problems representative of real-world scenarios.
IV-B4 GridMaze (Fig. 3(d))
@@
IV-C Experimental Results
In the Warehouse scenario (Tab. I), all methods except MRdRRT achieve a 100% success rate in the 2-robot scenario. However, the methods that do not leverage guidance begin to struggle with the 4-robot scenario and multiple cannot solve the 8-robot scenario. Of those, only ARC is able to solve up to the 32-robot scenario; however, with a success rate of only 26.6%. CDR-RRT and sRRT are able to solve up to the 8-robot scenario where the exponential size of the composite space leads to a significant decrease in performance. Workspace Guided-DaSH, like CDR-RRT, sees the benefit of using the skeleton representation to guide planning through narrow passages and the coordination provided by using the composite space of multiple robots. Additionally, by using the workspace to inform the planner when to move between different levels of composition, Workspace Guided-DaSH is able to focus planning within smaller spaces than CDR-RRT, making it able to plan for significantly larger robot teams, up to 128 robots, within the 600-second time limit. In this largest 128 robot scenario, the runtime of our MAPF heuristic was $16.7\pm 6.4$ seconds, demonstrating that it maintains scalability for large robot groups as well.
In the Tunnels environment (Tab. II), the size of the search space is increased significantly relative to the Warehouse environment since the robot now has six degrees of freedom. Composite RRT, MRdRRT, and ARC fail to solve the 2-robot scenario due to the complexity of the environment. sRRT performs better because of its ability to decompose the search space, but struggles with narrow passages and only one seed was able to solve the 4-robot scenario. Here, the narrow passage problem is exacerbated since the query requires the planner to find and traverse multiple narrow passages. By leveraging skeleton guidance, CDR-RRT and Workspace Guided-DaSH are able to find solutions for larger groups of robots. Searching the composite space hinders CDR-RRT’s performance, leaving it unable to solve the 64-robot scenario within the time limit. Workspace Guided-DaSH is able to efficiently find solutions for up to 64 robots within the time limit due to its ability to decompose the planning space.
@@ IV-C Experimental Results
In the Mining environment (Tab. III), Composite RRT and MRdRRT fail to solve the 2-robot scenario due to the large size of the search space and inter-robot conflicts. Similar to the Tunnels scenario, ARC’s performance is impacted by the environment complexity since it does not use topological guidance and it fails to solve the 2-robot scenario. sRRT struggles to solve the 2-robot scenario and cannot solve the 4-robot scenario. By leveraging topological guidance to limit exploration of the composite space, CDR-RRT is able to solve up to the 16-robot scenario. However, it is unable to solve the 32-robot scenario. Workspace Guided-DaSH, by considering the composite space only when necessary, is able to efficiently solve up to the 64-robot scenario.
@@
In the 3D GridMaze scenario (Tab. IV), as a result of the large size of the planning space as well as the difficulty posed by the long curved narrow passages, Composite RRT struggles with the 2-robot scenario and fails to solve the 4-robot scenario. MRdRRT fails to solve the 2-robot scenario. While CDR-RRT is able to solve the 4-robot scenario, the large size of the composite space prevents it from solving the 8-robot scenario within the time limit. sRRT and ARC benefit from searching a lower-dimensional space when robots are not in conflict, but fail to solve the 32-robot scenario where conflicts become prevalent due to the high level of congestion of the environment. Similar to sRRT, Workspace Guided-DaSH searches lower-dimensional spaces, however, our method also uses the workspace skeleton for conflict resolution. When a conflict is too difficult to be resolved quickly via random sampling, we impose constraints on the robots’ paths over the workspace skeleton which allows us to find conflict-free paths more efficiently. Thus, Workspace Guided-DaSH is able to find solutions up to 32 robots. Although this replanning mechanism sometimes results in slightly higher solution costs relative to other methods, a post-processing trajectory optimization step is generally used in practice to smooth paths.
IV-D Kinodynamic Experiments